# Gradient-based optimization of exact stochastic kinetic models

### v1

@@ A Reparameterized Pathwise Gradient Estimator
The inverse design objective is to maximize $J$ at fixed $\sigma$ and $K_{\text{tot}}$. The exact cycle current (net reaction counts divided by time) is not differentiable with respect to rate parameters. We instead estimate the current from simulated trajectories as
where $a_{i}^{\pm}(t_{n})$ are the forward and backward propensities and $\Delta t_{n}$ is the waiting time at step $n$. This time-weighted propensity average is differentiable and provides gradients with respect to all rate parameters. The entropy production estimate follows as $\hat{\sigma}=\hat{J}\cdot\mathcal{A}$. Since thermodynamic behavior depends on rate ratios $k_{i}^{+}/k_{i}^{-}$, the same physical driving can arise from many absolute rate values. We constrain all backward rates to a common value (still learnable given the fixed total budget), establishing a reference scale that removes this degeneracy and makes the optimized forward rates directly comparable across edges. Constraints are enforced through logarithmic penalties, ensuring scale-invariant treatment across the range of target values:
For the symmetric ring with fixed kinetic budget $K_{\text{tot}}$, the Pareto-optimal current at affinity $\mathcal{A}$ follows from steady-state analysis (Section S2):
The optimized cycle currents closely track the theoretical Pareto front across the full range of target dissipation values (Fig. 4b). Agreement extends from the near-equilibrium regime ($\sigma\ll 1$), where $J\propto\sqrt{\sigma}$ (dotted line) [28], to the far-from-equilibrium regime ($\sigma\gtrsim 1$), where current saturates toward its maximum value $J_{\text{max}}=K_{\text{tot}}/9$. To assess constraint satisfaction, we evaluate relative errors in the optimized current (compared to the theoretical optimum), entropy production rate, and total kinetic budget (compared to their target values). The kinetic budget constraint is satisfied to within fractions of 1% across all runs (Fig. 4c, green triangles; Table S1), ensuring valid comparison with the theoretical bound. Relative errors in $J$ and $\sigma$ remain small throughout the moderate-to-high dissipation regime but increase at the lowest entropy production targets, where the near-equilibrium current becomes vanishingly small and stochastic fluctuations dominate.
The optimized common backward rate $k^{-}$ and mean forward rate $\langle k^{+}\rangle$ closely follow theoretical predictions (Fig. 4d). The individual forward rates, instead, display substantial spread near equilibrium before converging to the fully symmetric allocation at high dissipation (Fig. 4d, light green symbols). This behavior is consistent with the close agreement between optimized currents and the Pareto bound: at small $\sigma$, the optimal rate configuration is only weakly determined at the level of individual edges. Many distinct allocations of the forward-rate budget yield indistinguishable $J(\sigma)$ once $\sigma$ and $K_{\text{tot}}$ are fixed, producing a broad family of near-degenerate solutions. Physically, although the smallest forward rate would nominally act as a bottleneck, transitions in the near-equilibrium regime are so infrequent that the delay from a single slow edge becomes negligible compared to the long waiting times throughout the cycle. Far from equilibrium, cycles complete rapidly enough that any rate asymmetry creates a persistent bottleneck, reducing the attainable current and driving the system toward symmetric rate allocation. Detailed optimization trajectories for each target $\sigma$ are provided in Figs. S10S19.
Discussion
@@
The ST-GS estimator introduces a biased approximation to gradient computation, controlled by a single temperature hyperparameter $\tau$. Crucially, this bias affects only the backward pass, while forward trajectories remain exact samples from the Chemical Master Equation throughout optimization. While theoretical characterization of the bias remains an open question, we find that the default value $\tau=1$ yields sufficiently accurate gradients for precise parameter recovery across diverse kinetic regimes and close agreement with analytical predictions in the thermodynamic optimization problem. This suggests the approach is robust for practical applications without requiring problem-specific tuning.
Differentiability does not, however, eliminate the intrinsic difficulty of inverse problems in stochastic kinetic systems. The loss landscapes exhibit the flat ridges and sloppy parameter directions characteristic of kinetic models, requiring careful optimizer configuration. We deliberately avoided problem-specific manipulations such as reparameterizations to better-conditioned variables or non-dimensionalization schemes that, while potentially easing optimization, would obscure the general applicability of the method. Full convergence was nonetheless achieved using modified Adam hyperparameters. Second-order methods that account for local curvature may offer further improvements but are more susceptible to noise in stochastic gradient estimates. The feasibility of gradient-based optimization demonstrated here may motivate systematic study of loss landscape geometry in kinetic models and the design of inference algorithms tailored to this problem structure, paralleling historical developments in neural network optimization.
A principal motivation for gradient-based approaches is that reverse-mode automatic differentiation scales independently of parameter dimensionality, in principle enabling efficient inference in large mechanistic models. We have deliberately restricted this study to low-dimensional systems where analytical solutions or ground-truth parameters permit rigorous validation. Extending to higher-dimensional systems is a natural direction but will require addressing the compounding challenges of optimization in increasingly degenerate landscapes.
A significant limitation of this work is the absence of experimental data. Demonstrating robust inference on real gene expression measurements—subject to technical noise, incomplete detection, and biological heterogeneity—will be necessary to establish the practical utility of this framework. Additionally, integration with variational inference methods such as ADVI could extend the approach to full posterior estimation, expanding the reach of Bayesian inference in biophysics and stochastic thermodynamics.
Finally, stochastic kinetic models extend well beyond the applications considered here, finding use in epidemiology, ecology, neuroscience, and the social sciences. The ability to efficiently differentiate through exact stochastic simulations provides a foundation for large-scale inference and principled investigation of optimality across the many domains governed by continuous-time Markov dynamics.
@@ Methods
Fully differentiable, hardware-accelerated simulations were coded using the JAX python library [7] and other libraries in the JAX ecosystem, in particular Equinox [20] as the simulation backbone. JAX provides automatic differentiation, just-in-time compilation and automatic vectorization that greatly improve computational performace of Monte Carlo gradient estimation. We provide a succinct overview of the main components of our model. A full mathematical account of simulation and optimization details is given in the Supplementary Information.
@@
We generate 25 parameter sets $\{k_{\text{on}},k_{\text{off}},k_{\text{tx}},k_{\text{mdeg}}\}$ chosen to span a range of biologically plausible kinetic regimes characteristic of bacterial gene expression. The off-rate is fixed at $k_{\text{off}}=5.0~\text{min}^{-1}$, corresponding to a mean ON-state duration ($1/k_{\text{off}}$) of approximately 12 seconds. This implies short transcriptional bursts, consistent with rapid transcription factor binding and unbinding dynamics. The mRNA degradation rate $k_{\text{mdeg}}$ ranges from 0.12 to 0.7 min-1, corresponding to half-lives of approximately 1–6 minutes. This range encompasses typical E. coli mRNA lifetimes, where the average half-life is 2–5 minutes. The transcription rate $k_{\text{tx}}$ spans 2.5–50 min-1. The lower end represents standard constitutive promoters, while rates approaching 50 min-1 (nearly one initiation per second) correspond to highly active promoters such as those driving ribosomal RNA synthesis or fully induced viral genes. These rates produce mean burst sizes ranging from 0.5 to 10 transcripts per burst, consistent with experimental measures in prokaryotic systems. The resulting mean mRNA copy numbers at steady state range from approximately 0.3 to 5 transcripts per cell, typical of low-to-moderate expression bacterial genes where the majority of transcripts are present at fewer than 10 copies per cell. This bursty, low-copy number regime represents a challenging test case for inference due to the inherent sparsity and stochasticity of the data.
@@
Soft bin assignments are obtained by normalizing per sample, $w_{ij}=k_{ij}/\sum_{l}k_{il}$. This normalization ensures conservation of probability mass for samples near the grid boundaries, without biasing the estimation in the bulk. We employ a bandwidth equal to the bin width, corresponding to $h=1.0$ in the case of integer data in this work. This ensures that one grid point contributes exactly zero mass to neighboring bins, recovering the exact probability mass function. Crucially, this also maintains non-zero gradients for the bin assignments, yielding a differentiable estimator that avoids introducing smoothing artifacts for integer data, in contrast e.g. to the more common Gaussian kernel.
@@ Acknowledgements.
The Python library written to conduct this study is available at https://github.com/fmottes/stochastix/. Code and data for reproducing the experiments in this paper is available at https://github.com/fmottes/stochastix-paper/.
@@
S2 Derivation of the Pareto-Optimal Current Bound
$K_{\text{tot}}=\sum_{i=1}^{3}(k_{i}^{+}+k_{i}^{-}).$
The second is a thermodynamic constraint fixing the cycle affinity,
In this symmetric regime, the constraints reduce to
$K_{\text{tot}}=3(k_{+}+k_{-}),\qquad\mathcal{A}=3\ln\frac{k_{+}}{k_{-}},$
which can be inverted to yield
$k_{+}+k_{-}=\frac{K_{\text{tot}}}{3},\qquad\frac{k_{+}}{k_{-}}=e^{\mathcal{A}/3}.$
$J=\pi_{i}k_{+}-\pi_{i+1}k_{-}=\frac{1}{3}(k_{+}-k_{-}).$
$J^{*}\approx\frac{K_{\text{tot}}}{54}\mathcal{A}.$
$J^{*}_{\max}=\frac{K_{\text{tot}}}{9},$
independent of the driving force.
S3 Variance-Reduced Gradient Estimation with Baseline Simulations
@@
Moment Matching
@@
Distribution Matching
@@
Three-state Non-equilibrium Cycle Model
@@
Additional Results. Table S1 reports the optimized current $J$, achieved entropy production $\sigma$, and total kinetic budget $K_{\text{tot}}$ for each target, along with relative errors compared to theoretical predictions. The kinetic budget constraint is satisfied to within 0.24% across all runs. Relative errors in current range from $6.9\%$ to $2.9\%$, with larger deviations at low $\sigma$ where the absolute current is small. Table S2 reports the optimized rate constants. At high dissipation ($\sigma\gtrsim 1$), the forward rates converge to the symmetric allocation $k^{+}_{ij}\approx K_{\text{tot}}/9\approx 0.333$. At low dissipation, individual forward rates show substantial heterogeneity while their mean remains close to theory. Fig. S7 shows the Pareto front, Fig. S8 the optimized rates, and Fig. S9 the absolute and relative errors. Figs. S10S19 show detailed optimization trajectories for each target $\sigma$.
@@
Figure S7: Pareto front for the three-state cycle model. Optimized steady-state current $J$ as a function of entropy production rate $\sigma$ for fixed kinetic budget $K_{\text{tot}}=1$. (Left) Linear scale. (Right) Log-log scale. Red curve: theoretical Pareto-optimal bound derived for symmetric rate allocation (Section S2). Blue points: optimized values.

### v2

@@ V Inverse design in nonequilibrium stochastic thermodynamics
The exact cycle current (net reaction counts divided by time) is not differentiable with respect to rate parameters. We instead estimate it from simulated trajectories as
$\hat{J}=\frac{1}{T}\sum_{n}\left[\frac{1}{L}\sum_{i=1}^{L}\left(a_{i}^{+}(t_{n})-a_{i}^{-}(t_{n})\right)\right]\Delta t_{n},$
where $a_{i}^{\pm}(t_{n})$ are the forward and backward propensities at bond $i$ and $\Delta t_{n}$ is the waiting time at step $n$. This time-weighted propensity average is differentiable and provides gradients with respect to all rate parameters, with the entropy production estimate following as $\hat{\sigma}=\hat{J}\,\mathcal{A}$. We optimize a maximum entropy production objective with a logarithmic penalty enforcing the forward-rate budget:
$\mathcal{L}=-\hat{\sigma}+\lambda_{K}\left[\log\frac{1/L\sum_{i}k_{i}^{+}}{\bar{k}^{+}}\right]^{2}.$
All optimizations start from the same random initial rates and are run to stable convergence using the modified Adam optimizer described in Methods.
For a ring with exclusion interactions and fixed mean forward rate $\bar{k}^{+}$, the steady-state current is maximized when the forward rates are uniform, $k_{i}^{+}\equiv\bar{k}^{+}$, yielding the fundamental diagram
$J^{*}(\rho)=(\bar{k}^{+}-k^{-})\rho(1-\rho),$
with a finite-size correction of L/(L1)L/(L-1) for a ring of $L$ sites. The parabolic density dependence reflects the competition between particle availability ($\rho$) and vacant sites for hopping ($1-\rho$), with maximum current at half-filling.
The optimized currents closely track the theoretical predictions across the full range of densities (Fig. 5b), following the finite-size prediction for $L=10$ and approaching the mean-field limit for $L=30$. Relative errors in optimized current remain within $\pm 3\%$ across all densities for both system sizes (Fig. 5c), demonstrating quantitative agreement with the analytical predictions. The optimization recovers both the optimal current magnitude and the optimal rate allocation strategy. Starting from random initializations with CV $\approx 0.3$$0.7$, the coefficient of variation of the optimized forward rates decreases monotonically during training and approaches zero within 100 updates, confirming convergence to the theoretically optimal uniform allocation (Fig. 5d). This holds reliably across both system sizes.
It is worth noting the scale of this optimization problem. For $L=30$ at half-filling, the underlying state space comprises $\binom{30}{15}\approx 10^{8}$ configurations, making master-equation-based optimization prohibitively expensive in general. Differentiable simulation sidesteps this problem by operating directly on sampled trajectories.
A complementary benchmark—characterizing the Pareto-optimal current-dissipation tradeoff in a three-state ring under simultaneous entropy production and kinetic budget constraints—is presented in Section S7, where the optimized solutions closely track the analytically derived bound across two orders of magnitude in dissipation.
VI Discussion
@@
The ST-GS estimator introduces a biased approximation to gradient computation, controlled by a single temperature hyperparameter $\tau$. Crucially, this bias affects only the backward pass: forward trajectories remain exact samples from the Chemical Master Equation throughout optimization. While a complete theoretical characterization of the bias remains an open question, the default value $\tau=1$ empirically yields sufficiently accurate gradients for effective parameter recovery across diverse kinetic regimes and close agreement with analytical predictions, without requiring problem-specific tuning.
With differentiable but exact SSA, any differentiable objective computable from stochastic trajectories becomes in principle accessible to efficient gradient-based optimization, including complex objectives defined over full probability distributions rather than summary statistics alone. The approach enables all of the optimizations demonstrated here to converge in under 15 minutes on a single GPU — often in a matter of minutes — with the potential of substantially accelerating iterative cycles of model refinement and hypothesis testing. Because computational cost scales with the number of sampled trajectories rather than with state-space size, the method extends naturally to systems with large configuration spaces that may prove challenging for alternative approaches.
These capabilities complement existing inference strategies, each of which involves different tradeoffs: explicit state-space representation for Finite State Projection [30], growing variance with trajectory length for likelihood ratio estimators [38] and poor scaling to high-dimensional parameter spaces for Approximate Bayesian Computation [48]. The ST-GS estimator trades unbiased gradients for exact forward dynamics, and through reverse-mode automatic differentiation provides gradient cost independent of parameter dimensionality — a distinct and practically useful point in this space of methods.
Our method also presents specific practical limitations. Backpropagating through long trajectories requires storing all intermediate states, imposing memory constraints that necessitate strategies such as the gradient-tracked and baseline trajectory split used here. Objectives that are not naturally differentiable — such as histograms and cycle currents — require carefully designed differentiable surrogates, and identifying such surrogates may not be straightforward for new problem classes. The ST-GS gradient bias, while empirically benign at $\tau=1$ across all applications considered here, lacks theoretical convergence guarantees, and no principled criterion exists for selecting the temperature in new settings. More broadly, differentiability does not eliminate the intrinsic difficulty of inverse problems in stochastic kinetic systems, whose loss landscapes typically exhibit flat ridges and sloppy parameter directions [18].
Several extensions are natural. Reverse-mode automatic differentiation scales independently of parameter dimensionality, enabling in principle inference in larger mechanistic models. Concurrent independent work applied the same estimator to optimization benchmarks in high-dimensional parameter spaces, confirming its scalability [50]. Time-dependent rate parameters would extend the framework to non-stationary processes, and integration with automatic differentiation variational inference techniques [25] could enable full posterior estimation. More broadly, stochastic kinetic models arise throughout epidemiology, ecology, neuroscience, and the social sciences. The ability to efficiently differentiate through exact stochastic simulations provides a foundation for scalable inference and rational design across these domains.
@@ Methods
End-to-end differentiable, hardware-accelerated simulations were coded using the JAX Python library [7] and other libraries in the JAX ecosystem, in particular Equinox [22] as the simulation backbone. JAX provides automatic differentiation, just-in-time compilation and automatic vectorization that greatly improve computational performance of Monte Carlo gradient estimation. We provide a succinct overview of the main components of our model. A full mathematical account of simulation and optimization details is given in the Supplementary Information.
@@
We generate 25 parameter sets $\{k_{\mathrm{on}},k_{\mathrm{off}},k_{\mathrm{tx}},k_{\mathrm{mdeg}}\}$ chosen to span a range of biologically plausible kinetic regimes characteristic of bacterial gene expression. The off-rate is fixed at $k_{\mathrm{off}}=5.0~\text{min}^{-1}$, corresponding to a mean ON-state duration ($1/k_{\mathrm{off}}$) of approximately 12 seconds. This implies short transcriptional bursts, consistent with rapid transcription factor binding and unbinding dynamics. The mRNA degradation rate $k_{\mathrm{mdeg}}$ ranges from 0.12 to 0.7 min-1, corresponding to half-lives of approximately 1–6 minutes. This range encompasses typical E. coli mRNA lifetimes, where the average half-life is 2–5 minutes. The transcription rate $k_{\mathrm{tx}}$ spans 2.5–50 min-1. The lower end represents standard constitutive promoters, while rates approaching 50 min-1 (nearly one initiation per second) correspond to highly active promoters such as those driving ribosomal RNA synthesis or fully induced viral genes. These rates produce mean burst sizes ranging from 0.5 to 10 transcripts per burst, consistent with experimental measures in prokaryotic systems. The resulting mean mRNA copy numbers at steady state range from approximately 0.3 to 5 transcripts per cell, typical of low-to-moderate expression bacterial genes where the majority of transcripts are present at fewer than 10 copies per cell. This bursty, low-copy number regime represents a challenging test case for inference due to the inherent sparsity and stochasticity of the data.
@@
Soft bin assignments are obtained by normalizing per sample, $w_{ij}=k_{ij}/\sum_{l}k_{il}$. This normalization ensures conservation of probability mass for samples near the grid boundaries, without biasing the estimation in the bulk. We employ a bandwidth slightly larger than the bin width, corresponding to $h=1.01$ for integer data. This preserves small but nonzero kernel overlaps at neighboring bin centers, ensuring that gradients remain well-defined everywhere while introducing negligible smoothing bias. The resulting estimator closely recovers the exact probability mass function for integer data, in contrast e.g. to the more common Gaussian kernel which introduces broader smoothing artifacts.
@@ Acknowledgements.
The Python library written to conduct this study is available at https://github.com/fmottes/stochastix/. Code and data for reproducing the experiments in this paper are available at https://github.com/fmottes/stochastix-paper/.
@@
S2 Variance-Reduced Gradient Estimation with Baseline Simulations
@@
S3 Moment Matching Loss – Additional Details
@@
S4 Distribution Matching Loss – Additional Details
@@
S5 Experimental Timecourse Data Fitting – Additional Details
Experiment Details. We infer all eight kinetic parameters of the four-state promoter model ($k_{12}$, $k_{21}$, $k_{23}$, $k_{32}$, $k_{34}$, $k_{43}$, $k_{\text{tx}}$, $k_{\text{mdeg}}$) from nuclear STL1 mRNA count distributions measured by smFISH at eight time points ($t=0,2,4,6,8,10,15,20$ min) during the rising phase of the osmotic stress response, using data from Li & Neuert [26]. Counts from three biological replicates were pooled at each time point. All trajectories were initialized in state $S_{1}$ with zero mRNA. We minimized the sum of per-time-point Kullback–Leibler divergences between observed and simulated RNA count distributions. Differentiable histograms were constructed using triangular kernel density estimation with bandwidth $h=1.01$ on a fixed grid of 11 points over $[0,10]$. Each optimization used 1024 gradient-tracked simulations combined with a variable number of baseline (forward-only) simulations per epoch. The baseline count was adjusted per time point so that the total pool matched the number of observed cells at that time point. Rate constants were parameterized as $k=e^{\phi}$ and optimization was performed over the log-transformed variables $\phi$, initialized uniformly at $\phi_{0}=-2$. Training ran for 100 epochs using standard Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and learning rate $\text{lr}=0.05$.
Additional Results. Table S1 reports the optimized rate constants. Fig. S7 shows the loss trajectories. Figs. S8S9 show the fitted distributions compared to the pooled target data alongside the initial (pre-optimization) model predictions. Figs. S10S11 compare the fitted model distributions against individual biological replicates at all eight time points, complementing the comparison shown in the main text (Fig. 4d).
Parameter Reaction Rate (min-1) $\log k$
$k_{12}$ $S_{1}\to S_{2}$ 0.237 1.438-1.438
$k_{21}$ $S_{2}\to S_{1}$ 0.068 2.691-2.691
$k_{23}$ $S_{2}\to S_{3}$ 0.296 1.217-1.217
$k_{32}$ $S_{3}\to S_{2}$ 0.047 3.062-3.062
$k_{34}$ $S_{3}\to S_{4}$ 0.384 0.957-0.957
$k_{43}$ $S_{4}\to S_{3}$ 0.042 3.169-3.169
$k_{\text{tx}}$ $S_{4}\to S_{4}+\text{RNA}$ 0.434 0.835-0.835
$k_{\text{mdeg}}$ $\text{RNA}\to\varnothing$ 0.095 2.354-2.354
Table S1: Optimized rate constants for the four-state promoter model.
Figure S7: smFISH fitting loss trajectories. KL divergence (training objective, left) and CDF $L_{1}$ distance (logged for reference, right) over 100 training epochs.
Figure S8: smFISH distribution fitting results (log scale). Comparison of RNA count distributions at all eight time points. Black: target distribution; gray dotted: initial model (pre-optimization); red: fitted model. The $y$-axis shows probability density on a logarithmic scale.
Figure S9: smFISH distribution fitting results (linear scale). Same data as Fig. S8 plotted with linear probability axes.
Figure S10: Fitted model vs. individual biological replicates (log scale). Comparison of RNA count distributions at all eight fitted time points. Colored circles: individual replicates (1–3); red circles: fitted model. The $y$-axis shows probability density on a logarithmic scale.
Figure S11: Fitted model vs. individual biological replicates (linear scale). Same data as Fig. S10 plotted with linear probability axes.
S6 ASEP Current Maximization – Additional Details
Experiment Details. We optimized the $L$ site-dependent forward hopping rates $\{k_{i}^{+}\}_{i=1}^{L}$ of a reversible ASEP ring to maximize steady-state current under a fixed mean forward rate budget $\bar{k}^{+}=\sum_{i}k_{i}^{+}/L=1$. The backward rate was fixed homogeneously at $k^{-}=0.01$ for all bonds. We swept nine particle counts for each system size ($L=10$: $N=1,\ldots,9$; $L=30$: $N=3,6,\ldots,27$), covering densities $\rho=0.1,0.2,\ldots,0.9$. Rather than directly maximizing current, we equivalently maximized entropy production $\hat{\sigma}=\hat{J}\,\mathcal{A}$, where $\hat{J}$ is the time-weighted propensity current estimate and $\mathcal{A}=\sum_{i}\log(k_{i}^{+}/k^{-})$ is the cycle affinity. The mean forward rate constraint was enforced via a logarithmic penalty with weight $\lambda=100$. Forward rates were parameterized as $k_{i}^{+}=e^{\phi_{i}}$ and optimized in log-space, initialized with log-normal disorder ($\phi_{i}\sim\mathcal{N}(0,0.5^{2})$, corresponding to CV $\approx 0.5$). Each optimization used batches of 8 trajectories ($L=10$) or 16 trajectories ($L=30$), each simulated for $T_{\text{sim}}=100$ time units. Training ran for 100 epochs using Adam with modified hyperparameters ($\beta_{1}=0.8$, $\beta_{2}=0.9$) and learning rate $\text{lr}=0.05$.
Additional Results. Figs. S12S20 show training traces and forward rate profiles (initial and optimized) for $L=10$ at all nine densities; Figs. S21S29 show the corresponding results for $L=30$.
S6.1 L=10: Convergence to the Finite-Size Current Optimum
The red dashed line in the training traces indicates the mean-field current $J_{\text{MF}}=(\bar{k}^{+}-k^{-})\rho(1-\rho)$. For $L=10$, the optimized current converges not to the mean-field prediction but to the finite-size corrected value $J^{*}=J_{\text{MF}}\cdot L/(L-1)$, as shown in the main text (Fig. 5b).
Figure S12: ASEP optimization details, $L=10$, $\rho=0.1$ ($N=1$). (Top) Training dynamics: loss, optimized approximate current $J$ (blue), exact discrete current (gray), and entropy production rate (green) over training iterations. The red dashed line indicates the mean-field current $J_{\text{MF}}$ (Bottom) Forward rate profiles across bonds at initialization (orange) and at convergence (blue); dashed line indicates the target mean rate $\bar{k}^{+}$.
Figure S13: ASEP optimization details, $L=10$, $\rho=0.2$ ($N=2$). Same format as Fig. S12.
Figure S14: ASEP optimization details, $L=10$, $\rho=0.3$ ($N=3$). Same format as Fig. S12.
Figure S15: ASEP optimization details, $L=10$, $\rho=0.4$ ($N=4$). Same format as Fig. S12.
Figure S16: ASEP optimization details, $L=10$, $\rho=0.5$ ($N=5$). Same format as Fig. S12.
Figure S17: ASEP optimization details, $L=10$, $\rho=0.6$ ($N=6$). Same format as Fig. S12.
Figure S18: ASEP optimization details, $L=10$, $\rho=0.7$ ($N=7$). Same format as Fig. S12.
Figure S19: ASEP optimization details, $L=10$, $\rho=0.8$ ($N=8$). Same format as Fig. S12.
Figure S20: ASEP optimization details, $L=10$, $\rho=0.9$ ($N=9$). Same format as Fig. S12.
S6.2 L=30: Approach to the Mean-Field Limit
For $L=30$, the finite-size correction $L/(L-1)\approx 1.03$ is small, and the optimized current closely approaches the mean-field prediction across all densities (Fig. 5b).
Figure S21: ASEP optimization details, $L=30$, $\rho=0.1$ ($N=3$). Same format as Fig. S12.
Figure S22: ASEP optimization details, $L=30$, $\rho=0.2$ ($N=6$). Same format as Fig. S12.
Figure S23: ASEP optimization details, $L=30$, $\rho=0.3$ ($N=9$). Same format as Fig. S12.
Figure S24: ASEP optimization details, $L=30$, $\rho=0.4$ ($N=12$). Same format as Fig. S12.
Figure S25: ASEP optimization details, $L=30$, $\rho=0.5$ ($N=15$). Same format as Fig. S12.
Figure S26: ASEP optimization details, $L=30$, $\rho=0.6$ ($N=18$). Same format as Fig. S12.
Figure S27: ASEP optimization details, $L=30$, $\rho=0.7$ ($N=21$). Same format as Fig. S12.
Figure S28: ASEP optimization details, $L=30$, $\rho=0.8$ ($N=24$). Same format as Fig. S12.
Figure S29: ASEP optimization details, $L=30$, $\rho=0.9$ ($N=27$). Same format as Fig. S12.
S7 Optimal Current-Dissipation Trade-Off in a Three-State Ring
In addition to the ASEP problem presented in the main text, we apply our framework to a complementary inverse design problem in stochastic thermodynamics: characterizing the Pareto front of optimal trade-offs between steady-state current and entropy production in a driven system with fixed kinetic resources. This problem combines core ingredients of non-equilibrium thermodynamics (cycle currents, thermodynamic affinities, dissipation constraints) in a minimal setting that admits analytical solution, providing a rigorous benchmark.
We study a three-state ring [19] where a particle transitions cyclically among states $S_{1}\to S_{2}\to S_{3}\to S_{1}$, with each edge supporting both forward and backward transitions (Fig. S30a):
Breaking detailed balance induces a net cycle current $J$ with entropy production rate $\sigma=J\mathcal{A}$, where $\mathcal{A}=\sum_{i}\log(k_{i}^{+}/k_{i}^{-})$ is the cycle affinity [42]. We impose a kinetic budget $K_{\text{tot}}=\sum_{i}(k_{i}^{+}+k_{i}^{-})$ bounding the system’s dynamical activity [29].
The inverse design objective is to maximize $J$ at fixed $\sigma$ and $K_{\text{tot}}$. As in the ASEP problem, we estimate the current from time-averaged propensities along simulated trajectories:
where $a_{i}^{\pm}(t_{n})$ are the forward and backward propensities and $\Delta t_{n}$ is the waiting time at step $n$. This time-weighted propensity average is differentiable and provides gradients with respect to all rate parameters. The entropy production estimate follows as $\hat{\sigma}=\hat{J}\,\mathcal{A}$. Since thermodynamic behavior depends on rate ratios $k_{i}^{+}/k_{i}^{-}$, the same physical driving can arise from many absolute rate values. We constrain all backward rates to a common value (still learnable given the fixed total budget), establishing a reference scale that removes this degeneracy and makes the optimized forward rates directly comparable across edges. Constraints are enforced through logarithmic penalties, ensuring scale-invariant treatment across the range of target values:
For the symmetric ring with fixed kinetic budget $K_{\text{tot}}$, the Pareto-optimal current at affinity $\mathcal{A}$ follows from steady-state analysis (Section S7.1):
The optimized cycle currents closely track the theoretical Pareto front across the full range of target dissipation values (Fig. S30b). Agreement extends from the near-equilibrium regime ($\sigma\ll 1$), where $J\propto\sqrt{\sigma}$ (dotted line) [34], to the far-from-equilibrium regime ($\sigma\gtrsim 1$), where current saturates toward its maximum value $J_{\text{max}}=K_{\text{tot}}/9$. To assess constraint satisfaction, we evaluate relative errors in the optimized current (compared to the theoretical optimum), entropy production rate, and total kinetic budget (compared to their target values). The kinetic budget constraint is satisfied to within fractions of 1% across all runs (Fig. S30c, green triangles; Table S2), ensuring valid comparison with the theoretical bound. Relative errors in $J$ and $\sigma$ remain small throughout the moderate-to-high dissipation regime but increase at the lowest entropy production targets, where the near-equilibrium current becomes vanishingly small and stochastic fluctuations dominate.
The optimized common backward rate $k^{-}$ and mean forward rate $\langle k^{+}\rangle$ closely follow theoretical predictions (Fig. S30d). The individual forward rates, instead, display substantial spread near equilibrium before converging to the fully symmetric allocation at high dissipation (Fig. S30d, light green symbols). This behavior is consistent with the close agreement between optimized currents and the Pareto bound: at small $\sigma$, the optimal rate configuration is only weakly determined at the level of individual edges. Many distinct allocations of the forward-rate budget yield indistinguishable $J(\sigma)$ once $\sigma$ and $K_{\text{tot}}$ are fixed, producing a broad family of near-degenerate solutions. Physically, although the smallest forward rate would nominally act as a bottleneck, transitions in the near-equilibrium regime are so infrequent that the delay from a single slow edge becomes negligible compared to the long waiting times throughout the cycle. Far from equilibrium, cycles complete rapidly enough that any rate asymmetry creates a persistent bottleneck, reducing the attainable current and driving the system toward symmetric rate allocation. Detailed optimization trajectories for each target $\sigma$ are provided in Figs. S34S43.
Figure S30: Optimal current-dissipation trade-off in a three-state ring. Optimization of cycle current $J$ under constraints on entropy production rate $\sigma$ and total kinetic budget $K_{\text{tot}}$. (a) Schematic of the three-state ring with forward ($k_{ij}^{+}$) and backward ($k_{ij}^{-}$) transition rates. (b) Optimized cycle current (blue circles) compared to the theoretical Pareto front (dashed line). Dotted line: asymptotic near-equilibrium scaling. Inset: Pareto front on linear axes. (c) Relative errors in $J$ (versus theory), $\sigma$, and $K_{\text{tot}}$ (versus target). Shaded band: $\pm 1\%$. (d) Optimized forward rates $k_{12}^{+}$, $k_{23}^{+}$, $k_{31}^{+}$ (light green symbols), their mean $\langle k^{+}\rangle$ (green diamonds), and backward rate $k^{-}$ (squares), compared to theoretical predictions (dashed and dotted lines).
S7.1 Derivation of the Pareto-Optimal Current Bound
$K_{\text{tot}}=\sum_{i=1}^{3}(k_{i}^{+}+k_{i}^{-}).$
The second is a thermodynamic constraint fixing the cycle affinity,
In this symmetric regime, the constraints reduce to
$K_{\text{tot}}=3(k_{+}+k_{-}),\qquad\mathcal{A}=3\ln\frac{k_{+}}{k_{-}},$
which can be inverted to yield
$k_{+}+k_{-}=\frac{K_{\text{tot}}}{3},\qquad\frac{k_{+}}{k_{-}}=e^{\mathcal{A}/3}.$
$J=\pi_{i}k_{+}-\pi_{i+1}k_{-}=\frac{1}{3}(k_{+}-k_{-}).$
$J^{*}\approx\frac{K_{\text{tot}}}{54}\mathcal{A}.$
$J^{*}_{\max}=\frac{K_{\text{tot}}}{9},$
independent of the driving force.
S7.2 Additional Details
@@
Additional Results. Table S2 reports the optimized current $J$, achieved entropy production $\sigma$, and total kinetic budget $K_{\text{tot}}$ for each target, along with relative errors compared to theoretical predictions. The kinetic budget constraint is satisfied to within 0.24% across all runs. Relative errors in current range from $6.9\%$ to $2.9\%$, with larger deviations at low $\sigma$ where the absolute current is small. Table S3 reports the optimized rate constants. At high dissipation ($\sigma\gtrsim 1$), the forward rates converge to the symmetric allocation $k^{+}_{ij}\approx K_{\text{tot}}/9\approx 0.333$. At low dissipation, individual forward rates show substantial heterogeneity while their mean remains close to theory. Fig. S31 shows the Pareto front, Fig. S32 the optimized rates, and Fig. S33 the absolute and relative errors. Figs. S34S43 show detailed optimization trajectories for each target $\sigma$.
@@
Figure S31: Pareto front for the three-state cycle model. Optimized steady-state current $J$ as a function of entropy production rate $\sigma$ for fixed kinetic budget $K_{\text{tot}}=1$. (Left) Linear scale. (Right) Log-log scale. Red curve: theoretical Pareto-optimal bound derived for symmetric rate allocation (Section S7.1). Blue points: optimized values.